From GlueXWiki
Jump to: navigation, search

February 14, 2013 FDC meeting


  1. Production Construction Tracking (Dave)
  2. Basis of Estimate for the FDC activities (Lubomir, Bill, Dave)
  3. Corrosion tests at MePhi latest results (Lubomir)
  4. Engineering (Bill)
  5. Electronics (Chris, Nick)
  6. Tests at EEL126 FDC E-log (Beni)
  7. Other


Participants: Eugene, Bill, Dave, Chris, Nick, Simon, Beni, and Lubomir.


- Dave for the spare parts: wire frame #3 (old) was tested without HV caps (580/-50nA at 2100/-500V) and now Anatoly is soldering them. Last wire frame (#32) was strung by Casey and Dave, Casey is gluing the wires now, tomorrow will come soldering and tension measurements. Don't need to convert type-2 to type-1 cathode frame (mill the O-ring groove) we have enough; the one that was deformed was used and looks OK. No need for extra spacer rings, the ones on the full-scale prototype can be used (Bill). Tina will come over the weekend to work on the last cathodes. Expect to finish with the spare part production in March.

- Lubomir for package #2: completely tested, oxygen went down to 170ppm. Started installing the grounding, 75% done. It is now on the mounting table and we don't have extra strong-back (used in 126) to put it on a table, or we can't use the holding brackets to hang it on the rails. The holding brackets will be used to test the deformation of the installation cart. So, next week we will move it on a granite table so that we can start working on the spare package.

Basis of Estimate for the FDC

- New version of the BOE is linked above. Most of the discussions were after the meeting with Eugene and on the next day.

Corrosion test at MePhi

- Already 18 days Viton and Viton+Apiezon samples are OK. Vlad mentioned that due to the condensation there is some water also directly in the cup with the samples, so the conditions are even more severe.


- Bill went step by step over the procedure (linked above) for bending/cutting the cooling tubes and installing the manifolds. Beni was concerned about the radiation hardness of the Tygon tubing that will be used for cooling. We expect to have 1kRad for 10 years and this is not a problem for the Tygon (Bill).

- Test of the installation cart deformation: we don't need the survey people, will just get the tool from Jim to measure the distance between the rails.

- 208V in 126 for the chiller will cost us $4.5k. At the same time Beni mentioned that there's 208V already and he's using it for one of the crates. He may need it also for another two crates when he loads them fully with electronics. Beni will check how much current is needed for that. Bill will check also what is the current needed for the chiller (it has breaker at 20A). Eugene insisted to consult also with Giles about the codes, if we can do it temporary by ourself.

Tests at EEL126

- Beni tested all the strip and wire channels (entries 104 and 105 in the FDC E-log). All channels are fine except two strip without signal, but could be electronics or cable connection. One wire channel has lower efficiency and also needs to be checked. Eugene suggested to measure the wire efficiencies at different thresholds and HV. Beni has some problems with the strip efficiencies, but is working on that.


- No meeting next week due to the collaboration meeting.